OpenStudy (anonymous):

(-4square root -5)(-2 Square root 3)

OpenStudy (anonymous):

multiply and simplify. assume that all variables represents nonnegative numbers. (-4 Square root -5)(-2 Square root 3)

OpenStudy (austinl):

\[-4 \sqrt{-5} \times -2 \sqrt{3}\]

OpenStudy (anonymous):

that is correct that Is the problem

OpenStudy (austinl):

Ok, so when we have a negative under the square root, what does that mean we will have in the final answer?

OpenStudy (anonymous):

a -

OpenStudy (austinl):

No, it means we will have an imaginary number. A non-real number.

OpenStudy (anonymous):

ok so no I am lost on this

OpenStudy (anonymous):

ok so now I am lost on this

OpenStudy (austinl):

Ok, the way this would work is you take the numbers outside the square root and multiply them. Can you do that, and what do you get?

OpenStudy (anonymous):

-8

OpenStudy (austinl):

-2*-4=8 Then you would multiply the numbers under the square roots.

OpenStudy (anonymous):

and inside the square root would be 15

OpenStudy (austinl):

\[8\sqrt{-15}\] But you can't have a negative under the square root so we pull it out and put an "i" next to the 8. \[8i\sqrt{15}\] This is your final answer.

OpenStudy (anonymous):

thank you so much for your help

OpenStudy (austinl):

No problem! Not gonna try and be rude here, but it is considered polite to give a medal to a helpful response.

OpenStudy (anonymous):

Their you go